OpenStudy (anonymous):

1/2a+2/3=-1/2a-3/4 solve the following equation.

OpenStudy (anonymous):

add 1/2 a to both sides

OpenStudy (anonymous):

do you want the answer or a way to solve it?

OpenStudy (anonymous):

both would supply.

OpenStudy (anonymous):

1/2a+2/3=-1/2a-3/4 if you added "1/2 a" left and right, what do you get ?

OpenStudy (anonymous):

1/2a+1/2a+2/3=-3/4

OpenStudy (anonymous):

correct you can combine the "a terms": \[\frac{ 1 }{ 2 }a+\frac{ 1 }{ 2 }a=a\]

OpenStudy (anonymous):

a+2/3=-3/4

OpenStudy (anonymous):

yes. now all that remains is to move 2/3 to the other side because the fractions do not have the same denominator, it is necessary to find a common denominator. it is best to find the least common factor

OpenStudy (anonymous):

\[a = -\frac{3}{4}-\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

so the answer it a=-3/2-2/3 ?

OpenStudy (anonymous):

*-3/4

OpenStudy (anonymous):

you can still simplify the right hand side -3/2 and -2/3 are both numerical terms, so it's possible to combine them.

OpenStudy (anonymous):

-3/4*

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } + \frac{2}{4}=\]it is possible to add fractions if you get them to the same denominator first

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

extending or reducing fraction changes the denominator, without changing the value

OpenStudy (texaschic101):

you can solve it that way...with fractions OR this way, without fractions... 1/2a + 2/3 = -1/2a - 3/4 (multiply the entire equation by the LCD, which is 12. 12(1/2)a + 12(2/3) = 12(-1/2)a - 12(3/4) 6a + 8 = -6a - 9 (add 6a to both sides) 6a + 6a + 8 = -9 (subtract 8 from both sides) 6a + 6a = -9 - 8 (combine like terms) 12a = - 17 (divide by 12) a = -17/12 or - 1 5/12 either way you choose to do it, the answers will be the same :)

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } =\frac{ 4 }{ 6 } \]\[=\frac{ 8 }{ 12 }\]

OpenStudy (anonymous):

now im confused.. which one is correct?

OpenStudy (anonymous):

both are correct. texaschic ones avoids dealing with fractions

OpenStudy (anonymous):

but which is in simplified form?

OpenStudy (anonymous):

we were left here: \[a = -\frac{3}{4}-\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

texaschic how did you find the LCD = 12?

OpenStudy (anonymous):

which gets the same end result: \[a = -\frac{9}{12}-\frac{ 8 }{ 12 }= -\frac{ 17 }{ 12 }\]

OpenStudy (texaschic101):

either way you choose to do the problem, the end result is the same

OpenStudy (anonymous):

woah. woah.. so a=-2/3-3/4 is a=−9/12−8/12=−17/12?

OpenStudy (texaschic101):

correct....when adding/subtracting fractions, they have to have the same denominator.

OpenStudy (texaschic101):

if hey don't, you have to make them by finding the LCD of the fractions

OpenStudy (texaschic101):

typo * they

OpenStudy (anonymous):

LCD.. how did they become this way? a=−9/12−8/12

OpenStudy (texaschic101):

ok...let me try to explain... -2/3 - 3/4 cannot be added together because they have different denominators (bottom numbers). So we have to make them the same. We have to find what number 3 and 4 both go into evenly. That would be 12. So we are going to make 12 the denominator. -2/3.....how many times does 3 go into 12 ? It goes in there 4 times, then take the 4 and multiply it by the numerator (2). That means 4 x 2 = 8. Put the 8 over the new denominator 12 giving us -8/12. -3/4....how many times does 4 go into 12 ? It goes in there 3 times. Now multiply 3 by the numerator (3) and you get 9. Put the 9 over the new denominator 12 giving us -9/12. Now we have : -8/12 - 9/12 (same denominator so we can add). -8/12 - 9/12 = -17/12 Understand ?

OpenStudy (anonymous):

omg xD you explain way better than our professor.. :"> thanks madam..

OpenStudy (texaschic101):

anytime...glad I could help :)